8. You buy stock and its price rises at a rate of five percent. Inflation for the same period rises at a rate of five percent. Before taxes you made
A. a nominal and real loss, but you pay taxes on the real loss.
B. a nominal and real gain, and you pay taxes on the nominal gain.
C. a nominal and real gain, but you pay taxes only on the real gain.
D. a nominal gain, but no real gain, so you pay no taxes on the capital gain.
E. a nominal gain, but no real gain, yet you pay taxes on the nominal gain.


9. Archie purchases some land for $30,000. He maintains it, but makes no improvements to it. One year later he sells it for $32,000. Betty puts $30,000 in a savings account that pays 6% interest. Archie has to pay a 50% capital gains tax, Betty is in the 30% tax bracket. The inflation rate was 2%. Who had the higher before-tax real gain and who had the higher after-tax real gain?
A. Archie had both the higher before-tax real gain and the higher after-tax real gain.
B. Archie had the higher before-tax real gain but Betty had the higher after-tax real gain.
C. Betty had the higher before-tax real gain but Archie had the higher after-tax real gain.
D. Betty had both the higher before-tax real gain and the higher after-tax real gain.
E. Betty had the higher before-tax real gain but they had the same after-tax real gain.
Answer #1

8> E. a nominal gain, but no real gain, yet you pay taxes on the nominal gain.

Reason

The tax you pay on capital gain is not inflation-adjusted. Thus, you pay tax based on the nominal value.

9> Before tax, nominal gain of Archie is 2000/30000=6.667%

Real Gain of Archie is 4.667%

Real Gain of Betty is 6%-2%=4%

After tax, real gain of Archie is 4.667(1-0.5)%=2.333%

After tax, real gain of Betty is 4(1-0.3)%=2.8%

So, B is correct